# Hardware test and validation of the angular droop control: Analysis and experiments

### v1

@@ Abstract
The angular droop control is a grid-forming control strategy that exploits the idea of power-to-angle droop to achieve exact frequency synchronization with no stringent separation between primary and secondary frequency control. In this work, we conduct hardware experiments in the Smart Energy System Control Laboratory at Karlsruhe Institute of Technology (KIT) to test and validate the angular droop control for low voltage power grids in two different test scenarios. First, we verify its grid-forming capabilities after a major event, e.g., following a blackout, demonstrated via power-to-angle droop behavior. For this, we propose two implementation schemes that rely either on direct or indirect actuation of the modulation signal and draw a comparison between them. Second, we investigate the plug-and-play capabilities, i.e., local stability and power sharing for a two-converter system and provide suitable tuning for the control gains. Our experimental findings illustrate the usefulness of hardware test and validation for DC/AC converter control, the practical challenges entailed and the proposed remedies.
@@ I Introduction
Power grids are facing a rapid transition from fossil fuel towards an increasing share of renewable energy resources. This profound change is characterized by the integration of converter-based generation [1]. In particular, the high penetration of power electronics alters the power system dynamics, governed thus far by rotating synchronous machines underpinning the legacy grid [2]. Therefore, the control of DC/AC converters lies at the forefront of this transition to ensure the stability of power systems [3, 4, 5].
For a given control scheme, the sole deployment of numerical simulations for validation is unsatisfactory due to the discrepancy between real-world setup and simplified settings adopted in numerical case studies such as anomalous and unmodeled dynamics, erroneous model parameters and unknown disturbances affecting the grid [5]. In fact, conducting hardware experiments in controlled laboratory environments offers the possibility to integrate hardware prototypes and is thus a significant step towards field application of new control paradigms. For this, a power-hardware-in-the-loop testbed for intelligent operation and control of low-inertia power systems is proposed in [6]. Therein, the facility is specifically designed for experimentally testing and validating control schemes for low-inertia power systems. In the same vein, the Smart Energy System Control Laboratory at KIT is introduced in [7]. It consists in a fully-automated and user-oriented research infrastructure for controlling and operating smart energy systems thanks to its high level of automation and the capacity to fully function in a grid-decoupled way. This allows to test, study and evaluate control algorithms in a safe environment.
@@
In this context, hardware-in-the-loop tests have been conducted on several controllers to study their closed-loop capabilities under realistic operating conditions. One of the most well-studied controllers in converter-based power grid is the frequency droop control initially proposed in [8]. Inspired by the dynamics governing synchronous machines and their analogy to coupled oscillator dynamics [9], frequency droop control is a grid-forming method extensively studied both in theory and practice. An experimental validation of the droop control strategy can be found in [10] for a single and two droop-controlled converters. The dynamics of synchronous machines remain a source of inspiration for a multitude of converter control strategies that emulate the behavior of synchronous machines. One particular controller that relies on exact model matching of high-order dynamics of three-phase synchronous machines with three-phase balanced and averaged DC/AC converters derived from first-order principles is the matching control introduced in [11, 12]. The particularity of this controller relies on easily measured DC-side voltage representing an indicator of power imbalance in the grid without prior assumptions on quasi-stationary steady state and the operation on phasor quantities. An experimental setup consisting of two voltage source converters for the matching control is proposed in [13] and performed on a general-purpose test bench. Oscillator-based control schemes such as the virtual oscillator control [14] rely on emulating the dynamics of weakly coupled nonlinear oscillators. Compared to frequency droop control, which is only well-defined in the vicinity of a sinusoidal steady state, virtual oscillator control enables interconnected converters to stabilize synchronous sinusoidal waveforms starting from arbitrary initial condition. The virtual oscillator control is validated experimentally in [15] within a laboratory hardware prototype to demonstrate the validity of the design approach. Even though the virtual oscillator control has provable droop properties [16], the control tuning remains a difficult task due to a lack of intuition on the physical interpretation of the gains. It was also not possible to track active and reactive power setpoints in its original formulation in [14]. These limitations have motivated a variant of virtual oscillator control suggested in [17] that allows active and reactive power to be dispatched, hence the name dispatchable Virtual Oscillator Control or d-VOC. The d-VOC consists of a combination of a synchronizing feedback term, together with a decentralized magnitude control law. It allows for global stabilization of the angles and voltage magnitudes at their desired values, corresponding to a pre-specified solution of the AC power-flow equations. The controller exhibits droop behavior around the standard operating point, which makes it backward compatible with existing power system components. Experimental validation of the dispatchable virtual oscillator on a testbed are found in [18].
The angular droop control trades linearly active power with angle deviation at steady state and thus achieves exact frequency regulation. Compared to frequency droop control, the angular droop controller exploits the idea of power-to-angle droop to achieve frequency synchronization with zero stationary frequency error and thus merging primary with secondary frequency control [19]. The angular droop control is shown to be inverse optimal stabilizing for the angle dynamics under mild assumptions [20, 21]. Thereby, the converter phase angles are stabilized to an induced steady-state, where the convergence rate is traded with the allowed control effort. The optimality of the angular droop control brings about inherent desirable gain margins analogous to linear quadratic regulators and showcases the utility of inverse optimal control theory in networked settings [22]. The angular droop control has been studied for constrained DC/AC converter angles and overall power generation [23], for discrete-time systems [24] and numerically tested on several power system benchmarks [25, 26, 27, 20]. In the present work, we study the angular droop control in a laboratory environment provided by the Smart Energy System Control Laboratory at KIT from the lenses of hardware experimentation. First, we demonstrate that a detailed high-order model including a boost converter can be accounted for even if the theoretical analysis in [20, 21] relies on a simplified DC/AC converter model description. For this, we actuate the modulation signal representing the main input to the high-dimensional DC/AC converter. Second, we test the grid-forming capabilities of the angular droop control, namely the ability to form sinusoidal voltage waveforms at a desired AC frequency and amplitude. This is exemplified through the black start capabilities following a major event, e.g., a blackout as well as the ability to withstand load disturbances, i.e., upon a sudden increase/decrease in the load power. Third, motivated by the wide use of cascaded inner loops for the control of DC/AC converters, we propose an indirect implementation scheme for the angular droop control and draw a comparison to the direct scheme acting on the modulation signal. Fourth, we validate frequency synchronization and the power sharing capabilities of the angular droop control in dependence of the control gains shown in [20]. Therein, we provide useful insights into the tuning of the gain from a practitioner point of view. Finally, we highlight the practical challenges encountered within the different experiments and provide tractable solutions to circumvent them.
The paper is organized as follows: Section II introduces the angular droop control and discusses its properties. A detailed high-order modeling of the converters and their controls are presented in Section III. Section IV describes the experimental setup. Sections V and VI summarize the results of the test scenarios and Section VII concludes the paper.
@@
Notation: Define $\mathbf{I}=\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$ and $\mathbf{J}=\begin{bmatrix}0&-1\\ 1&0\end{bmatrix}.$ Let $\mathbb{T}^{n}$ denote the $n-$th dimensional torus. Consider a network described by a connected graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\Xi)$, consisting of $|\mathcal{V}|=n$ nodes representing DC/AC converter buses and $|\mathcal{E}|=m$ edges modeling purely inductive transmission lines (i.e., with zero conductances) with susceptance $b_{kj}>0,\;(k,j)\in\mathcal{E}$ collected in the diagonal matrix $\Xi=\mathrm{diag}(b_{kj}),\;(k,j)\in\mathcal{E}$. The topology of the graph $\mathcal{G}$ is described by the incidence matrix $\mathcal{B}\in^{n\times m}$. We denote by $\mathcal{N}_{k}$ the neighbor set of converter $k$. Let $x$ denote an AC quantity in $abc-$frame and $x_{dq}:=\mathcal{P}(\theta_{dq})\,x$ denote its transformation in $dq-$frame [2] following a Park transformation $\mathcal{P}(\theta_{dq})$ with angle $\theta_{dq}(t):=\theta_{k}(t)$.
@@ II The angular droop control
II-A Control law
Consider a network of DC/AC converters, each represented by a voltage phasor. Hereby all the phasors are modeled with a constant magnitude (e.g. one per unit), and the converter’s angle dynamics are assumed to be controllable. We assume quasi-stationary steady state operation, i.e., the frequency is around a nominal steady state $\omega^{*}$ [2]. Overall the network dynamics can be represented by,
@@ II-A Control law
where $\hat{u}(\theta)=[\hat{u}_{1}(\theta),\dots,\hat{u}_{n}(\theta)]^{\top}\in\mathbb{R}^{n}$ is the main control input, $\theta=[\theta_{1},\dots,\theta_{n}]^{\top}\in\mathbb{R}^{n}$ is the vector of phase angles of the DC/AC converters and $\theta_{0}=[\theta_{0,1},\dots,\theta_{0,n}]^{\top}\in\mathbb{R}^{n}$ is the initial angle vector. By discarding the internal converter dynamics, we design a mathematical tractable approach to find an optimal controller in a concise and closed-form based on inverse optimality as shown in [20]. In Section III and IV, we demonstrate through hardware experimentation that a detailed high-order converter model can be accounted for.
Next, we denote the vector of nominal phase angles, rotating at nominal frequency $\omega^{*}>0$ by $\theta^{*}(t)=\omega^{*}\mathds{1}_{n}+\theta^{*}_{0}\in^{n}$, where $\theta^{*}_{0}=[\theta^{*}_{0,1},\dots,\theta^{*}_{0,n}]\in^{n}$ is the initial nominal angle vector. Let $P(\theta)=[P_{1}(\theta),\dots,P_{e,n}(\theta)]^{\top}\in^{n}$ be the vector of active power injected into the network and $P^{*}=[P^{*}_{1}(\theta),\dots,P^{*}_{e,n}(\theta)]^{\top}\in^{n}$ is the vector of nominal active power. The angular droop control is given by,
@@
with $R=\text{diag}(\alpha_{1},\dots,\alpha_{n})>0,\,\Gamma=\text{diag}(\gamma_{1},\dots,\gamma_{n})>0$. To design the angular droop control (2), we assume available synchrophasor measurements with respect to a global frame of reference. Our assumption is reasonable since phasor measurements are available at the Smart Energy System Control Laboratory at KIT [7]. See also Section IV. In summary, the closed-loop angle dynamics are given by,
$\dot{\theta}=-\frac{1}{2}R^{-1}\big{(}\Gamma\,(\theta-\theta^{*})+P(\theta)-P^{*})+\omega^{*}\mathds{1}_{n}.$ (3)
II-B Properties: Stability and optimality
@@
[…]
@@
Under Assumption 1, the angular droop control is the optimal stabilizing solution of (II-B) in a neighborhood of the induced steady state angle $\tilde{\theta}^{s}$ that satisfies
@@
where $P^{s}(\tilde{\theta}^{s})=[P_{1}^{s}(\tilde{\theta}^{s}),\dots,P_{n}^{s}(\tilde{\theta}^{s})]^{\top}$ is vector of the induced steady state active power [20] . Note that (5) describes the steady state as a power balance between the active power and angle deviation from the nominal value. It coincides with the steady state resulting from letting the running cost in (II-B) go asymptotically to zero, i.e.,
@@
The angle droop behavior (5) leads to zero stationary frequency error. In fact, by taking the time derivative of (5), we arrive at $\dot{\tilde{\theta}}^{s}=0$ and therefore $\dot{\theta}^{s}_{k}=\omega^{*}$. Thus, the steady state frequency error is zero and we conclude that:
  • The angular droop control (5) is able to guarantee both primary and secondary frequency control at once.
  • Local asymptotic stability of the induced steady state angle $\theta^{s}$ implies local frequency synchronization of all converters at the nominal value $\omega^{*}$.
II-C Control tuning
After selecting the gain matrices $R=\alpha\,\mathbf{I}_{n}$ and $\Gamma=\gamma\,\mathbf{I}_{n}$ with $\alpha,\gamma>0$, i.e., the control gains are uniform across all the converters, observe that:
  • The gain $\gamma>0$ defines the power-to-angle droop behavior between the power and angle deviation at steady state.
  • Both the gains $\gamma$ and $\alpha>0$ affect the rate of change of frequency or RoCoF given by $\dot{\omega}_{k}$. This can be seen as follows. By taking the time derivative of (3), we obtain,
    $\ddot{\theta}_{k}=-\frac{1}{2\alpha}\big{(}\gamma(\dot{\theta}_{k}-\dot{\theta}^{*}_{k})+\dot{P}_{k}(\theta)\big{)},\quad k=1\dots n$
    where $\dot{P}_{k}(\theta)=\frac{\mathrm{d}P_{k}(\theta)}{\mathrm{d}t}$. By letting $\omega_{k}:=\dot{\theta}_{k}$ and $\omega^{*}_{k}:=\dot{\theta}^{*}_{k}$, we have
    $\dot{\omega}_{k}=-\frac{1}{2\alpha}\big{(}\gamma\,(\omega_{k}-\omega^{*})+\dot{P}_{k}(\theta)\big{)},\quad k=1\dots n\,.$ (6)
    Therefore, a sudden change in active power corresponds to a sudden RoCoF $\dot{\omega}_{k}$ that occurs during the transients, while the frequency error remains zero at steady state. In (6), the RoCoF depends on both gains $\alpha>0$ and $\gamma>0$.
III High-order converter modeling and control
With some abuse of notation, we omit throughout subsequent sections the subscript $k$ to denote a quantity $z_{k}$ of the $k-$th converter.
III-A Modeling
In this section, we relax our modeling assumptions from Section II towards a hardware experimentation in two directions:
  • First, we include a DC power supply and a DC/DC converter put in series behind the DC/AC converter and suggest a control approach for the boost converter relying on vector control consisting of nested voltage and current control loops.
III-A1 Boost converter
Figure 1: A schematic representation of the boost converter under study
$L_{b}\dot{I}_{b}$ $=-R_{b}\,I_{b}+V_{b}-V_{c}$ (7)
$C_{dc}\dot{V}_{dc}$ $=-G_{dc}V_{dc}+I_{dc},\quad I_{dc}=\frac{V_{b}I_{b}}{V_{dc}}\,.$
$V_{c}=(1-d)\cdot V_{dc}\in.$ (8)
III-A2 DC/AC converter
Figure 2: A three-phase DC/AC converter model under study
In the remainder, we consider a three-phase averaged and balanced DC/AC converter as shown in Fig. 2 and described in $abc-$frame [2, 28],
$L\,\dot{i}$ $=-R\,i+\frac{1}{2}\overline{u}\,\,V_{dc}-v,$ (9)
$C\,\dot{v}$ $=-G\,v+i-i_{o},$
where $V_{dc}$ denotes the DC side voltage from the boost converter model (7). On the AC side, let $i\in^{3}$ denote the inductance current and $v\in^{3}$ the output voltage. The filter resistance and inductance are specified by $R>0$ and $L>0$, respectively. The capacitor $C>0$ is set in parallel with the load conductance $G>0$ to ground. The DC/AC converter is connected to the AC network, where $i_{o}\in$ is the output current flowing into the network. Note that the pulse width modulation signal $\overline{u}\in[-1,1]$ relates to the converter duty cycle $d_{AC}\in[0,1]$ via
$d_{AC}=\frac{1}{2}+\frac{\overline{u}}{2},$ (10)
where $\overline{u}$ represents the main input to the DC/AC converter.
III-B Control schemes
In the remainder, we present control schemes based on vector control for the boost converter and angular droop control for the DC/AC converter.
III-B1 Boost converter
The cascaded control architecture of the boost converter exploits the differential equations (7) and is summarized in Fig. 3. In particular, an outer loop regulates the DC capacitor voltage $V_{dc}$ at a nominal value $V_{dc}^{*}>V_{b}>0$ by specifying a reference current $I^{d}_{b}\in$ given explicitly by,
$d=1-\frac{V_{b}-V^{d}_{l}}{V_{dc}}.$ (13)
III-B2 Control of DC/AC converter
We distinguish two approaches to implement the angular droop control (2) following direct or indirect control of the modulation signal $\overline{u}$ in (9).
Direct implementation
First, we define the active power $P=v^{\top}i_{o}$, and the nominal steady state $P^{*}=v^{*\top}i_{o}^{*}$ and propose the following control scheme
where $10<A<$ 1 is the amplitude of the modulation signal $\overline{u}$. Fig. 4 depicts a summarizing block diagram of a single DC/AC converter whose system dynamics are given by (9), set in closed-loop with the angular droop control (2). Note that in (14), the angular droop control increments the converter’s state with a virtual angle. In this scheme, the modulation signal $\bar{u}$ has the virtual angle $\theta$ given by (14).
Figure 4: Implementation of the angular droop for the DC/AC converter via direct control of the modulation signal $\overline{u}$.
Indirect Implementation
In this section, we propose an indirect implementation of the angular droop control that relies on vector control. In particular, the indirect implementation entails inner voltage and current control loop according to a cascaded architecture.
After a Park transformation $\mathcal{P}(\theta_{dq})$ with angle $\theta_{dq}(t):=\theta(t)$ and given the reference voltage $v^{d}$ in $abc$-frame [2],
$i_{dq}^{d}$ $=Y\,v_{dq}+i_{o,dq}-k_{VP}(v_{dq}-v_{dq}^{d})$ (16)
$-k_{VI}\int_{0}^{t}\big{(}v_{dq}(\tau)-v_{dq}^{d}\big{)}\mathrm{d}\tau,$
where $k_{VP},k_{VI}~{}>0$ are the control gains and $Y=G+C\mathbf{J}\omega^{*}$. To track the reference current $i_{dq}^{d}$ in (16), we design an inner current loop based on proportional and integral control using the switching voltage $v^{d}_{m}=\frac{1}{2}\overline{u}_{dq}\,V_{dc}$ as follows
$v_{m}^{d}=Z\,i_{dq}+v_{dq}-k_{IP}(i_{dq}-i^{d}_{dq})-k_{II}\int_{0}^{t}(i_{dq}(\tau)-i_{dq}^{d})\,\mathrm{d}\tau,$
where $k_{IP},k_{II}>0$ are control gains and $Z=R+L\,\mathbf{J}\,\omega^{*}$. By applying the inverse Park transformation $\mathcal{P}^{-1}(\theta_{dq})$, we recover the modulation input $\overline{u}$ in $abc-$frame as follows,
It is noteworthy that, the pairs $(k_{VP},k_{VI})$ and $(k_{IP},k_{II})$ are again chosen to guarantee time-scale separation, where the current control loop is faster than the voltage control loop. Fig. 5 summarizes the overall scheme of the indirect implementation based on the vector control.
Figure 5: Indirect cascaded control of the angular droop after $dq-$transformation of the DC/AC converter.
IV Experimental Setup
The following hardware experiments are conducted at the Smart Energy System Control Laboratory at KIT. For a comprehensive description of the overall lab equipment, see [7].
IV-A Hardware components
The hardware testbed consists of a programmable DC/AC converter system, a resistive load and transmission line replicas.
IV-A1 Programmable DC/AC converter system

### v2

@@ Abstract
We present a hardware-based validation of angular droop control for grid-forming DC/AC converters, a control strategy that establishes active power-to-angle droop. Angular droop control enables exact frequency regulation at steady state, thereby combining primary and secondary control into a single layer. We provide traceable analysis and suggest solutions to the main implementation challenges with angular droop control, specifically addressing the challenges concerning discretization and clock drift in hardware experiments. This is illustrated in two different scenarios. Experimental results from the single converter to load scenario demonstrate black start capability and power-to-angle droop behavior for two different implementation schemes. A multi-converter setup validates frequency synchronization and power-sharing properties, proving the ancillary services that angular droop control provides in the real-world experimental setup.
@@ I Introduction
Power grids are facing a rapid transition from fossil fuel towards an increasing share of renewable energy resources. This profound change is characterized by the integration of converter-based generation [28]. In particular, the high penetration of power electronics alters the power system dynamics, governed thus far by rotating synchronous machines underpinning the legacy grid [34]. Therefore, the control of DC/AC converters lies at the forefront of this transition to ensure power system stability [9].
@@
Typical control approaches for grid-forming DC/AC converters are inspired by the dynamics governing synchronous machines and their analogy to coupled oscillator dynamics [31]. For example, frequency droop control is based on the active power to frequency droop, which is inherent to synchronous machines and enables the synchronization between different generators. This grid-forming method is extensively studied both in theory and practice. An experimental validation of the frequency droop control strategy for different applications can be found in [5]. The dynamics of synchronous machines remain a source of inspiration for a multitude of converter control strategies that emulate their behavior. One particular controller that relies on exact model matching of high-order dynamics of synchronous machines is the matching control introduced in [16, 3]. This controller relies on easily measured DC-side voltage to play the role of an indicator of power imbalance in the grid. Oscillator-based control schemes such as the Virtual Oscillator Control [15] rely on emulating the dynamics of weakly coupled nonlinear oscillators. Compared to frequency droop control, which is only well-defined in the vicinity of a sinusoidal steady state, virtual oscillator control enables interconnected converters to stabilize synchronous sinusoidal waveforms starting from arbitrary initial condition. The virtual oscillator control is validated experimentally in [14] within a laboratory hardware prototype to demonstrate the validity of the design approach. Even though the virtual oscillator control has provable active power to frequency droop properties [32], the control tuning remains a difficult task due to a lack of intuition on the physical interpretation of the gains. It was also not possible to track active and reactive power setpoints in its original formulation in [15]. These limitations have motivated a variant of Virtual Oscillator Control suggested in [6] that allows active and reactive power to be dispatched, hence the name dispatchable Virtual Oscillator Control whose experimental validation is conducted in [30]. The authors in [33] design an angle based modification for a frequency droop controller to dampen power oscillations in a multiple converter setup.
All the control approaches discussed above have in common that they are based on the principle of active power to frequency droop. In contrast, the angular droop control studied in this paper establishes a linear relationship between active power and angle deviation, rather than frequency. This achieves exact frequency regulation at steady state. As a consequence, it merges primary with secondary frequency control [25] and no additional control layer, i.e., secondary control is necessary. During transients, compared to frequency droop control, the angular droop acts on the rate of change of frequency to counterbalance the rate of change of power which anticipates a change in the power balance itself. Therefore angular droop control reacts faster to load disturbances. Additionally, the angular droop control is shown to be inverse optimal stabilizing for the angle dynamics [17, 22]. The optimality of the angular droop control brings about inherent desirable gain margins analogous to linear quadratic regulators and showcases the utility of inverse optimal control theory in networked settings [18]. Angular droop control has been tested in simulations on numerous power system benchmarks. In [25] the operation of a simulated microgrid consisting of angular droop controlled DC/AC converters is presented and the zero frequency deviation property at steady state is shown in comparison to frequency droop control. Kolluri et. al [23] examine the power sharing properties of angular droop control and validate their results using a simulation study. Xu et. al [37] deploy a simulation study of an islanded microgrid to validate an angular droop based method for cost minimization and consensus active power sharing between distributed generation units. In [17] it is proven that angular droop is an inverse optimal locally stabilizing control law for a multi-converter system, which is supported by numerical simulations. Those results are extended [20] to account for input and output constraints posed by real-world applications and the tuning of discrete-time implementations of angular droop control is discussed in [19]. Both studies support their results using numerical simulations.
The sole deployment of numerical simulations for validation is unsatisfactory due to the discrepancy between real-world setups and simplified settings adopted in numerical case studies such as unmodeled dynamics and erroneous or unknown model parameters affecting the system.
This work demonstrates the grid-forming properties of angular droop control in a controlled experimental setting [36], with particular emphasis on the hardware-based validation. The main contributions are as follows:
  • This paper presents the first hardware-based implementation of angular droop control. The challenges of opting for angle- instead of frequency-based control are addressed in a real hardware experimental setup.
  • Grid-forming properties such as black start, robustness to load changes, and zero steady state frequency deviation property of angular droop control are validated using the single converter-to-load scenario. The control law is reformulated to suit hardware realization, and two implementation schemes are compared to show compatibility with different inner control architectures.
  • Frequency synchronization and power-sharing capabilities are proven in the two-converter scenario, forming the basis for generalization to $n$-converter setups. Clock drift issues are analyzed and mitigated via master clock distribution. Further, practical guidance on control tuning for angular droop is provided.
This experimental validation represents a meaningful contribution toward the practical deployment of angular droop control, helping to bridge the gap between theoretical development and real-world implementation.
The paper is structured as follows. Section II first briefly introduces angular droop control and discusses the difference to frequency droop control, then details the experimental environment. Thereafter, in Section III the results of the single converter to load scenario are discussed along with the challenge of discretizing the angle dynamics. Section IV is concerned with the two-converter to load scenario and the challenges connected to synchronization and power-sharing. Lastly, Section V summarizes the findings and concludes the paper.
@@
Notation: Define $\mathbf{I}=\left[\begin{smallmatrix}1&0\\ 0&1\end{smallmatrix}\right]$ and $\mathbf{J}=\left[\begin{smallmatrix}0&-1\\ 1&0\end{smallmatrix}\right].$ Let $x$ denote an AC quantity in abcabc-frame and $x_{dq}:=\mathcal{P}(\theta_{dq})\,x$ denote its transformation in $dq$-frame [24] following a Park transformation $\mathcal{P}(\theta_{dq})$ with angle $\theta_{dq}(t):=\theta_{k}(t)$, where $\theta_{k}$ is the angle of converter $k$. Consider a network described by a connected graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\Xi)$, consisting of $|\mathcal{V}|=n$ nodes representing DC/AC converter buses and $|\mathcal{E}|=m$ edges modeling purely inductive transmission lines with susceptance $b_{kj}>0,\;(k,j)\in\mathcal{E}$ collected in the diagonal matrix $\Xi=\mathrm{diag}(b_{kj}),\;(k,j)\in\mathcal{E}$. The topology of the graph $\mathcal{G}$ is described by the incidence matrix $\mathcal{B}\in^{n\times m}$.
@@ II The angular droop control
II-A Control scheme
Consider a network of DC/AC converters, each represented by a voltage phasor. Hereby all the phasors are modeled with a constant magnitude (e.g. one per unit), and the converter’s angle dynamics are assumed to be controllable. Overall the network dynamics can be represented by,
@@ II-A Control scheme
where $\hat{u}(\theta)=[\hat{u}_{1}(\theta),\dots,\hat{u}_{n}(\theta)]^{\top}\in\mathbb{R}^{n}$ is the main control input, $\theta=[\theta_{1},\dots,\theta_{n}]^{\top}\in\mathbb{R}^{n}$ is the vector of phase angles of the DC/AC converters, $\theta_{0}=[\theta_{0,1},\dots,\theta_{0,n}]^{\top}\in\mathbb{R}^{n}$ is the initial angle vector and $\omega^{*}$ is the nominal angular frequency. The angular droop control is given by [17, 22],
@@
with $R=\text{diag}(\alpha_{1},\dots,\alpha_{n})>0,\,\Gamma=\text{diag}(\gamma_{1},\dots,\gamma_{n})>0$. Further, $P(\theta)=[P_{1}(\theta),\dots,P_{n}(\theta)]^{\top}\in\mathbb{R}^{n}$ denotes the active power vector, and $\theta^{*}=[\theta^{*}_{1},\dots,\theta^{*}_{n}]^{\top}\in\mathbb{R}^{n}$ and $P^{*}=[P^{*}_{1}(\theta),\dots,P^{*}_{n}(\theta)]^{\top}\in\mathbb{R}^{n}$ the respective reference values. Note that for the angular droop control design (2), we assume that synchrophasor measurements of the angle, with respect to a global frame of reference, are available to each converter. This is a reasonable scenario for a future power grid, as phasor measurement unit (PMU) installation is becoming increasingly widespread[11, 35, 10]. In summary, the closed-loop angle dynamics are given by,
$\dot{\theta}=-\frac{1}{2\alpha}\big(\gamma\,(\theta-\theta^{*})+P(\theta)-P^{*})+\omega^{*}\mathds{1}_{n},$ (3)
where the gain matrices $R=\alpha\,\mathbf{I}_{n}$ and $\Gamma=\gamma\,\mathbf{I}_{n}$ with $\alpha,\gamma>0$, i.e., the control gains are uniform across all the converters. Observe that:
  • The gain $\gamma>0$ defines the power-to-angle droop behavior between the power and angle deviation at steady state characterized by
    $\Gamma\,(\theta^{s}-\theta^{*})=P^{*}-P^{s}(\theta^{s}),$ (4)
    where $\theta^{s}\in^{n}$ and $P^{s}(\theta^{s})$ are the vectors of induced steady state angles and powers, respectively.
  • Both the gains $\gamma$ and $\alpha$ affect the rate of change of frequency or RoCoF given by $\dot{\omega}$. This can be seen by taking the time derivative of (3) as follows
    $\ddot{\theta}=-\frac{1}{2\alpha}\big(\gamma(\dot{\theta}-\dot{\theta}^{*})+\dot{P}(\theta)\big),$ (5)
    where $\dot{P}(\theta)=\frac{\mathrm{d}P(\theta)}{\mathrm{d}t}$. By letting $\omega:=\dot{\theta}$ and $\omega^{*}:=\dot{\theta}^{*}$, we have
    $\dot{\omega}=-\frac{1}{2\alpha}\big(\gamma\,(\omega-\omega^{*})+\dot{P}(\theta)\big).$ (6)
    Therefore, a sudden change in active power corresponds to a sudden RoCoF $\dot{\omega}$ that occurs during the transients, while the frequency error remains zero at steady state. Observe that the RoCoF depends on both gains $\alpha$ and $\gamma$.
Stability and optimality [17]
@@
$\text{s.t. }\dot{\tilde{\theta}}=\hat{u}(\tilde{\theta}),\quad\tilde{\theta}(0)=\tilde{\theta}_{0}.$ (7)
@@
Under Assumption 1, the angular droop control is the optimal stabilizing solution of (II-A) in a neighborhood of the induced steady state angle $\tilde{\theta}^{s}$ that satisfies
@@
where $P^{s}(\tilde{\theta}^{s})=[P_{1}^{s}(\tilde{\theta}^{s}),\dots,P_{n}^{s}(\tilde{\theta}^{s})]^{\top}$ is vector of the induced steady state active power [17] . Note that (4) describes the steady state as a power balance between the active power and angle deviation from the nominal value. It coincides with the steady state resulting from letting the running cost in (II-A) go asymptotically to zero, i.e.,
@@
II-B Angular vs. frequency droop – Distinction of properties
In this section, we highlight the merits of the angular droop control compared to frequency droop control in terms of frequency support and steady state behavior. In particular, we provide a comprehensive analysis that demonstrates why angular droop advances the state-of-the art.
Inspired by the dynamics governing synchronous machines and their analogy to coupled oscillator dynamics [31], frequency droop control is a grid-forming method that obeys the following closed-loop dynamics:
$\dot{\omega}$ $=-\frac{1}{2\alpha}\big(\gamma\,(\omega-\omega^{*})+P(\theta)-P^{*})\,.$ (9)
For easiness of comparison, we hereby denote the virtual inertia $M=2\,\alpha\,\mathbf{I}_{n}$ and the damping coefficient $D=\gamma\,\mathbf{I}_{n}$.
Linking active power to angle instead of frequency yields altogether different behavior during both transient and steady state, by comparing the frequency droop dynamics (9) to angular droop dynamics (3), we observe that:
Frequency support
For the angular droop control, the rate of change of frequency or RoCoF is induced by the rate of change of power in (6) and not the power imbalance itself as in (9). This means that the angular droop acts on the rate of change of frequency to counterbalance the rate of change of power which anticipates a change in the power balance itself. This is analogue to a derivative control that extrapolates the current slope of the error and generates one large corrective effort immediately after a load change in order to begin eliminating the error as quickly as possible. A loop with derivative control recovers quicker from a disturbance with less deviation than a loop with only a proportional term as in frequency droop control.
Steady state behavior
At steady state, the frequency droop behavior established by (9) is given by
$\gamma\,(\omega^{s}-\omega^{*})=P^{*}-P(\theta^{s}).$ (10)
Here, the induced steady state frequency $\omega^{s}$ deviates from the nominal frequency $\omega^{*}$ and depends directly on the magnitude of the disturbance affecting the power network. Therefore a secondary control is necessary to bring the steady state frequency back to nominal. On the other hand, the steady state frequency induced by angular droop control, see (4), settles to its nominal value at steady state in disregard of the disturbance amplitude implying zero frequency error and thus merging primary and secondary control [25]. Please refer to Fig. 9 in Sec. III-C for the experimental results on this comparison. This saves additional control effort while achieving the same result.
II-C Experimental setup
The hardware testbed consists of programmable DC/AC converter systems, a resistive load and transmission line replicas. Relevant parameters of the setup are summarized in Table I. Details about the model under consideration are presented in Appendix Hardware test and validation of the angular droop control: Analysis and experiments. For further information on the microgrid laboratory refer to [36].
Refer to captioncontrol unitPower Rack: containshalf-bridge modulesfor DC/DC andDC/AC converterAC output filter and DC-leg inductanceDC power supply
Figure 1: Overview of a programmable DC/AC converter system for the hardware experiments.